# LoFi: Scalable Local Image Reconstruction with Implicit Neural Representation

### v1

@@ V-B Reconstruction analysis and interpretation
Figure 7: The learned patch geometry for different pixels. As expected, larger $T$ results in more accurate localization of the relevant information.
@@
As discussed in Section III-C, LoFi adaptively estimates patch geometry for each pixel based on local information. Figure 7 illustrates the learned patch geometry across different pixels for the image denoising task, where the estimated positions of relevant features align with intuitive expectations. This ability to pinpoint features contributing to each pixel’s estimation can be valuable for more insightful analysis and interpretation. Additionally, the figure shows that more iterations (larger values of $T$) yield more precise localization of relevant information, enhancing the effectiveness of feature selection.
V-C Learning the imaging geometry with INR-LoFi
In this section, we study the performance of different models on a toy experiment where we have changes in the coordinate system. We consider recovering an image from its transpose; a simple task where we only have changes in the coordinate system while the image content remains unchanged. We train LoFi and CNNs on 29900 training samples from CelebA-HQ (Karras et al., 2017) dataset. Figure 8 showcases the performance of different models on both in-distribution (CelebA-HQ test samples) and OOD data (LSUN-bedroom (Yu et al., 2015)). Despite the simplicity of the task and access to a large-scale dataset, CNNs show poor performance as they can only learn the changes in the image content while neglecting the changes in the coordinate system. Unlike CNNs, our proposed INR-LoFi can conveniently learn these changes in the coordinate system resulting in significantly better reconstructions.
V-D Computational efficiency
As discussed in Section III-E, LoFi can operate on machines with extremely small GPUs by reducing the pixel batch size in both training and inference. We train all models with the same object batch size 64 on a dataset with 30000 training samples from CelebA-HQ dataset. Table II presents a comparative analysis of memory usage and training time (per epoch) across different models and image resolutions. This analysis highlights that, while CNN memory requirements scale with image resolution, LoFi’s memory usage remains nearly constant, requiring only 3GB for resolutions above $512\times 512$—over 20 times less than the typical memory needed for CNNs.
To address the significant memory demands of CNNs for large images, one common workaround is to train on smaller image patches. However, this approach is suboptimal due to the manual selection of patch size and the need to mitigate blocking artifacts at inference. Moreover, this patch-based training strategy is only effective for local tasks, such as image denoising, where the relevant information can be found within the corresponding input patch. For non-local tasks, like those discussed in Section V-C, this approach breaks down because the required information often lies outside the corresponding patch, making accurate reconstruction almost impossible.
TABLE II: Model comparison for memory usage and training time on one epoch.
Params $128\times 128$ $256\times 256$ $512\times 512$ $1024\times 1024$
DnCNN (Zhang et al., 2017a) 3M 23GB / 200s 78GB / 1680s $>80$GB $>80$GB
IRCNN (Zhang et al., 2017b) 3M 16GB / 120s 55GB / 500s $>80$GB $>80$GB
U-Net (Ronneberger et al., 2015) 8M 6GB / 40s 16GB / 100s 60GB / 380s $>80$GB
DRU-Net (Zhang et al., 2021) 8M 7GB / 60s 22GB / 220s 79GB / 800s $>80$GB
LoFi (with filter; ours) 4M 3GB / 80s 3GB / 100s 5GB / 140s 15GB / 300s
LoFi (no filter; ours) 3M 2GB / 60s 2GB / 80s 2GB / 120s 3GB / 260s
@@ V-E Low-resolution priors for high-resolution image reconstruction with LoFi-ADMM
Image in-painting. We train LoFi denoiser on 29900 samples from the CelebA-HQ (Karras et al., 2017) in low resolution $128\times 128$ where we consider zero-mean Gaussian noise ${\mathbf{n}}$ with $\sigma=0.15$ and normalize images between 0 and 1. In LoFi-ADMM, we set $\alpha=0.05$ and execute it for 90 iterations, the hyper-parameters we found suitable to ensure convergence. Now, we run LoFi-ADMM to solve the image-inpainting problem at a higher resolution $512\times 512$ where $p=30\%$ of the image pixels are randomly masked. Figure 9 showcases the performance of the LoFi-ADMM on both in-distribution and OOD data. This experiment shows that once LoFi denoiser is trained, it can be used for solving general inverse problems at any higher resolution with a strong generalization.
Figure 9: LoFi-ADMM performance on image in-painting ($p=30\%$) for in-distribution and OOD data in resolution $512\times 512$ when LoFi denoiser is trained on CelebA-HQ samples in low resolution $128\times 128$.
Radio interferometry. We next apply LoFi-ADMM to radio interferometric imaging, please refer to Section S-I-F in the supplementary material for further information regarding the forward model. We use non-uniform fast Fourier transform (NUFFT) to simulate radio interferometric measurements from $128\times 128$ images using an upsampling factor of 2 and Kaiser-Bessel kernels with a size of 6 × 6 pixels; we use the simulated uv-coverage of the MeerKAT telescope and the measurements with complex Gaussian noise at the SNR of 30dB. We train LoFi denoiser on 13000 images of simulated galaxies created from the IllustrisTNG simulations (Nelson et al., 2019). Figure 10 illustrates the performance of the LoFi-ADMM at original and higher resolutions. This experiment shows that LoFi-ADMM can conveniently exploit the measurements obtained from higher-resolution images for more accurate reconstructions with more details.
@@ VI Limitations and conclusion
We introduced LoFi, a scalable local image reconstruction framework that stands out for its strong generalization on OOD data and resolution agnostic memory usage during training. Notably, our approach achieves strong performance on local inverse problems such as LDCT and image denoising, all while maintaining a significantly smaller memory footprint compared to standard CNNs. The strong inductive bias of LoFi allows us to conveniently train it on tiny datasets with few training samples without overfitting or needing regularization like early stopping. Our experiments demonstrate the effectiveness of LoFi as a denoising prior within the PnP framework, showcasing its ability to reconstruct images in arbitrary resolutions for tasks such as image in-painting and radio interferometric imaging.
Although LoFi’s pixel-based pipeline significantly reduces memory requirements—particularly during training, thus enabling high-dimensional reconstructions on smaller GPUs—it comes with trade-offs. Like INRs, the computational cost of LoFi in inference increases quadratically with image resolution. Recent studies (He and Jin, 2023, 2024) have proposed strategies to enhance the efficiency of continuous image representation in INRs by increasing shared computations across pixels, thereby reducing computational complexity. Implementing similar strategies in LoFi could potentially reduce computations during inference, which we leave as a direction for future work.
@@
S-I Network architecture and training details
@@
S-I-A Low-dose computed tomography
@@
S-I-B Image denoising
@@
S-I-C Dark matter mapping
For this experiment, we used the same network architecture as the experiment of LDCT for both LoFi and baselines. Regarding dark matter mapping simulations, we used the $\kappa$TNG convergence maps [Osato et al., 2021], generated from the widely used IllustrisTNG hydrodynamical simulation suite [Pillepich et al., 2018, Springel et al., 2018, Naiman et al., 2018, Nelson et al., 2018, Marinacci et al., 2018], to build the train and test datasets for the models presented in Figure 6. The $\kappa$TNG dataset consists of 10 0001000010\,00010 000 realizations of $5\times 5\text{deg}^{2}$ convergence maps for each of the $40$ different source redshifts with $2.60<z_{s}<$ 2.6. For the sake of simplicity, we have selected a single redshift slice, the $20^{\text{th}}$ plane, corresponding to $z_{s}=0.858$ for our experiments.
From the simulated convergence fields, $\kappa$, we have constructed the observed shear fields using Equation 5. We follow [Price et al., 2020, §5.2.1] for the noise simulation. The shear noise ${\mathbf{n}}$ for each component of the shear and for each pixel $i$ is simulated as ${\mathbf{n}}_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$. The per-pixel standard deviation can be computed as follows
$\sigma_{i}=\frac{\sigma_{e}}{\sqrt{(\theta^{2}/n_{\text{grid}}^{2})\times n_{\text{gal}}}},$ (S1)
@@
S-I-D Training details
LoFi is implemented in PyTorch on a machine with an Nvidia A100 GPU with 80GB memory. LoFi and baselines were trained for 200 epochs with $\ell_{1}$ loss using Adam optimizer [Kingma and Ba, 2014] with learning rate $10^{-4}$ and object batch size 64. In the case of LoFi, for each mini-batch of random objects, we performed optimization on a random mini-batch of 512 pixels 2 times.
S-I-E Plug-and-Play ADMM
The PnP algorithm [Venkatakrishnan et al., 2013] is developed based on the alternating direction method of multipliers (ADMM) [Boyd et al., 2011], where new variables ${\mathbf{u}}$ and ${\mathbf{v}}$ are introduced to decouple the data fidelity and regularization terms in (1) as follows,
@@
$-\dfrac{1}{2\eta}\|{\mathbf{u}}\|_{2}^{2}\bigg{\}},$ (S2)
@@
${\mathbf{f}}_{k}=h({\mathbf{v}}_{k-1}-{\mathbf{u}}_{k-1};\alpha),$ (S3)
${\mathbf{v}}_{k}=\text{prox}_{R}({\mathbf{f}}_{k}-{\mathbf{u}}_{k-1};\eta),$ (S4)
${\mathbf{u}}_{k}={\mathbf{u}}_{k-1}+({\mathbf{f}}_{k}-{\mathbf{v}}_{k}),$ (S5)
@@
$h({\mathbf{z}};\alpha)\triangleq({\mathbf{A}}^{*}{\mathbf{A}}+\alpha)^{-1}({\mathbf{A}}^{*}{\mathbf{q}}+\alpha{\mathbf{z}}),$ (S6)
$\text{prox}_{R}({\mathbf{z}};\eta)\triangleq\operatorname*{argmin}_{\mathbf{f}}\dfrac{1}{2\eta}\|{\mathbf{f}}-{\mathbf{z}}\|_{2}^{2}+R({\mathbf{f}}),$ (S7)
@@
where ${\mathbf{A}}^{*}$ is the adjoint forward operator. The proximal operator in (S7) can be interpreted as the denoiser of ${\mathbf{z}}$ with image prior $R(\cdot)$ and AWGN variance $\eta$. The key idea of the PnP algorithm is to employ a powerful image denoiser, like a pre-trained CNN, for (S7).
S-I-F Radio interferometric imaging
@@
Radio interferometry is an imaging technique where we acquire samples of the Fourier transform of the image along curves [Thompson et al., 2017, Pratley et al., 2017, Liaudat et al., 2023]. In radio interferometric imaging, aperture synthesis techniques are used to acquire specific Fourier measurements giving an incomplete coverage of the Fourier plane. The inverse problem consists of recovering the entire intensity image from Fourier measurements. In this case, the forward operator in a simplified form writes
${\mathbf{A}}=\mathbf{M}_{\text{RI}}\mathbf{F}$ (S8)
@@
where $\mathbf{M}_{\text{RI}}$ is a Fourier mask indicating the Fourier coverage of the observations. We have omitted gridding and degridding operations in between other corrections from the forward operator for the sake of simplicity; a more detailed description of the operator can be found in  [Pratley et al., 2017, §3.2].
@@
Figure S1: Training dataset used for image denoising experiment in resolution $512\times 512$.
S-II Futher analysis and additional experiments
@@
In this section, we provide further analysis of LoFi architecture and additional experiments.
S-II-A High-dimensional image denoising
Figure S2 illustrates the reconstructed images by LoFi on image denoising task ($\sigma=0.2$) in resolution $1024\times 1024$ which confirms the expressivity of the proposed framework even on very large images.
@@
Figure S2: LoFi reconstructions on in-distribution and OOD data for image denoising ($\sigma=0.2$) in resolution $1024\times 1024$.
S-II-B Equivariance to shifts and rotations
@@
In many imaging and pattern recognition tasks, it is desirable to work with functions that are invariant or equivariant to shifts and rotations. Building such invariances into CNNs, typically via group representation theory [Cohen and Welling, 2016a, Bronstein et al., 2021] improves generalization both in and out of distribution [Cohen and Welling, 2016b, Veeling et al., 2018].
@@
It is easy to see that LoFi is shift equivariant by design. Indeed, the patch at position ${\mathbf{x}}+\Delta{\mathbf{x}}$ in an image shifted by $\Delta{\mathbf{x}}$ is exactly the same as the patch at position ${\mathbf{x}}$ in a non-shifted image; the $\text{NN}_{\theta}$ (MultiMLP) thus receives the same input. This stands in contrast to most mainstream multi-scale CNNs which are surprisingly sensitive even to 1-pixel shifts, although equivariance can be restored by a careful design of pooling and downsampling layers [Chaman and Dokmanić, 2021b, a].
@@
Rotation equivariance is precluded for square patches, except for angles that are multiples of $\frac{\pi}{2}$. With circular patches shown in Figure 2, when the image is rotated the $\text{NN}_{\theta}$ receives the same input only in a different order up to interpolation error. Rotation equivariance could be implemented by applying $\text{NN}_{\theta}$ to rotation invariant features [Lowe, 1999, Puny et al., 2021, Sannai et al., 2021, Zhao and Singer, 2014] extracted from the circular patch or by averaging the output over rotations. Even without this, however, LoFi is approximately rotation equivariant because natural images are approximately locally rotation invariant and the patch geometry is initialized as circular.
@@
Now, we assess LoFi performance on the image denoising task when the input image is shifted or rotated. As shown in Figure S3, we horizontally shifted the input noisy image for 15 pixels. As expected, LoFi’s reconstruction is shifted for 15 pixels and exhibits no degradation in the reconstruction quality. This experiment confirms that LoFi is a truly shift equivariant network.
@@
Figure S4 demonstrates the LoFi performance when the input noisy image undergoes a 90-degree rotation. As expected, we observe minimal degradation in reconstruction quality which shows LoFi’s strong robustness under rotation. Note that, we did not use any data augmentation to make the model robust under translation and rotation during training.
@@
Figure S3: Shift equivariance analysis for LoFi architecture; LoFi exhibits no degradation in reconstruction quality when the input noisy image is shifted for 15 pixels.
@@
Figure S4: Rotation equivariance analysis for LoFi architecture; LoFi exhibits minimal degradation in reconstruction quality when the input noisy image undergoes a 90-degree rotation.
S-II-C Learned filter
In this section, we analyze the learned filter ${\mathbf{H}}$ in (15) to understand how it processes noisy images. As shown in Figure S5, the learned filter remains local in data space after training which prevents $\text{NN}_{\theta}$ from overfitting the large-scale features.
Figure S5: The learned filter ${\mathbf{H}}$ in (15); as expected, it has a small support in data space which prevents the $\text{NN}_{\theta}$ from overfitting large-scale features.

### v2

@@ V-C Reconstruction analysis and interpretation
Figure 9: The learned patch geometry for different pixels. As expected, larger $T$ results in more accurate localization of the relevant information. The red dots indicate the query pixels, while the blue dots represent the estimated positions of relevant features to be extracted from the noisy image.
@@
As discussed in Section III-C, LoFi adaptively estimates patch geometry for each pixel based on local information. Figure 9 illustrates the learned patch geometry across different pixels for image denoising, where the estimated positions of relevant features align with intuitive expectations. This ability to pinpoint features contributing to each pixel’s estimation can be valuable for interpretation. Additionally, the figure shows that more iterations (larger values of $T$) result in a more precise localization of relevant information.
V-D Learning the imaging geometry with INR-LoFi
In this section, we study the performance of different models on a toy experiment where we have changes in the coordinate system. We consider recovering an image from its transpose; a simple task where we only have changes in the coordinate system while the image content remains unchanged. We train LoFi and CNNs on 29900 training samples from CelebA-HQ [84] dataset. Figure 10 showcases the performance of different models on both in-distribution (CelebA-HQ test samples) and OOD data (LSUN-bedroom [85]). Despite the simplicity of the task and access to a large-scale dataset, CNNs show poor performance as they can only learn the changes in the image content while neglecting the changes in the coordinate system. Unlike CNNs, our proposed INR-LoFi can conveniently learn these changes in the coordinate system resulting in significantly better reconstructions.
@@ V-E Low-resolution priors for high-resolution image reconstruction with LoFi-ADMM
Image inpainting
We train LoFi denoiser on 29900 samples from the CelebA-HQ [84] in low resolution $128\times 128$ where we consider zero-mean Gaussian noise ${\mathbf{n}}$ with $\sigma=0.15$ and normalize images between 0 and 1. In LoFi-ADMM, we set $\alpha=0.05$ and execute it for 90 iterations, the hyper-parameters we found suitable to ensure convergence. Now, we run LoFi-ADMM to solve the image-inpainting problem at a higher resolution $512\times 512$ where $p=30\%$ of the image pixels are randomly masked. Figure 11 showcases the performance of the LoFi-ADMM on both in-distribution and OOD data. This experiment shows that once LoFi denoiser is trained, it can be used for solving general inverse problems at any higher resolution with a strong generalization.
Figure 11: LoFi-ADMM performance on image in-painting ($p=30\%$) for in-distribution and OOD data at resolution $512\times 512$ when LoFi denoiser is trained on CelebA-HQ samples in low resolution $128\times 128$.
Radio interferometry
We next apply LoFi-ADMM to radio interferometric imaging, please refer to Section VII-F in the supplementary material for further information regarding the forward model. We use non-uniform fast Fourier transform (NUFFT) to simulate radio interferometric measurements from $128\times 128$ images using an upsampling factor of 2 and Kaiser–Bessel kernels with a size of 6 $\times$ 6 pixels; we use the simulated uv-coverage of the MeerKAT telescope and the measurements with complex Gaussian noise at the SNR of 30dB. We train LoFi denoiser on 13000 images of simulated galaxies created from the IllustrisTNG simulations [86]. Figure 12 illustrates the performance of the LoFi-ADMM at original and higher resolutions. This experiment shows that LoFi-ADMM can conveniently exploit the measurements obtained from higher-resolution images for more accurate reconstructions with more details.
@@ VI Limitations and conclusion
We introduced LoFi, a scalable local image reconstruction framework that stands out for its strong generalization on OOD data and resolution agnostic memory usage during training. LoFi achieves strong performance on local inverse problems such as LDCT and image denoising, all while maintaining a significantly smaller memory footprint compared to standard CNNs and vision transformers. The favorable inductive bias of LoFi allows us to train it on very small datasets without overfitting or explicit regularization. Our experiments further demonstrate the effectiveness of LoFi as a denoising prior within the PnP framework, showcasing its ability to reconstruct images at arbitrary resolutions for tasks such as image in-painting and radio interferometric imaging.
Although LoFi’s pixel-based pipeline significantly reduces memory requirements during training, it comes with trade-offs. Like INRs, the computational cost of LoFi in inference increases quadratically with image resolution as shown in Figure 8(b). Recent studies [87, 88] have proposed strategies to enhance the efficiency of continuous image representation in INRs by increasing shared computations across pixels, thereby reducing computational complexity. Implementing similar strategies in LoFi could potentially reduce computations during inference, which we leave as a direction for future work.
@@
VII Network architecture and training details
@@
VII-A Low-dose computed tomography
@@
Restormer: Transformer blocks = [4, 6, 6, 8], attention heads = [1, 2, 4, 8] and number of channels = [16, 32, 64, 128].
SwinIR: Patch embedding dimension = 120, patch size = 1, window size = 8, attention heads = [6,6,6,6,6,6].
VII-B Image denoising
@@
VII-C Dark matter mapping
For this experiment, we used the same network architecture as the experiment of LDCT for both LoFi and baselines. Regarding dark matter mapping simulations, we used the $\kappa$TNG convergence maps [89], generated from the widely used IllustrisTNG hydrodynamical simulation suite [90, 91, 92, 93, 94], to build the train and test datasets for the models presented in Figure 7. The $\kappa$TNG dataset consists of 10 0001000010\,00010 000 realizations of $5\times 5\text{deg}^{2}$ convergence maps for each of the $40$ different source redshifts with $2.60<z_{s}<$ 2.6. For the sake of simplicity, we have selected a single redshift slice, the $20^{\text{th}}$ plane, corresponding to $z_{s}=0.858$ for our experiments.
From the simulated convergence fields, $\kappa$, we have constructed the observed shear fields using Equation 5. We follow [95, §5.2.1] for the noise simulation. The shear noise ${\mathbf{n}}$ for each component of the shear and for each pixel $i$ is simulated as ${\mathbf{n}}_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$. The per-pixel standard deviation can be computed as follows
$\sigma_{i}=\frac{\sigma_{e}}{\sqrt{(\theta^{2}/n_{\text{grid}}^{2})\times n_{\text{gal}}}},$ (19)
@@ VII-C Dark matter mapping
VII-D Training details
LoFi is implemented in PyTorch on a machine with an Nvidia A100 GPU with 80GB memory. LoFi and baselines were trained for 200 epochs with $\ell_{1}$ loss using Adam optimizer [96] with learning rate $10^{-4}$ and object batch size 64. In the case of LoFi, for each mini-batch of random objects, we performed optimization on a random mini-batch of 512 pixels 2 times.
VII-E Plug-and-Play ADMM
In this section, we briefly review plug-and-play (PnP) framework for image reconstruction. PnP algorithms achieved SOTA performance across a variety of image reconstruction tasks including image super-resolution and deblurring [97, 15], computed tomography (CT) [98] and magnetic resonance imaging (MRI) [99, 100]; see also the recent review [101]. In PnP, a pre-trained Gaussian denoiser, typically a CNN, is used as a prior. Gaussian denoisers are often cheaper in both than generative priors [102, 103] in both training and inference. Training generative models also requires large-scale datasets and substantial computational resources. Recently, the authors of [104] leveraged diffusion models [103, 105] as a pre-trained denoiser in PnP framework and achieved excellent performance for various image restoration tasks.
Originally, PnP [78] is developed based on the alternating direction method of multipliers (ADMM) [106], where new variables ${\mathbf{u}}$ and ${\mathbf{v}}$ are introduced to decouple the data fidelity and regularization terms in (1) as follows,
@@ VII-E Plug-and-Play ADMM
$-\dfrac{1}{2\eta}\|{\mathbf{u}}\|_{2}^{2}\bigg{\}},$ (20)
@@
${\mathbf{f}}_{k}=h({\mathbf{v}}_{k-1}-{\mathbf{u}}_{k-1};\alpha),$ (21)
${\mathbf{v}}_{k}=\text{prox}_{R}({\mathbf{f}}_{k}-{\mathbf{u}}_{k-1};\eta),$ (22)
${\mathbf{u}}_{k}={\mathbf{u}}_{k-1}+({\mathbf{f}}_{k}-{\mathbf{v}}_{k}),$ (23)
@@
$h({\mathbf{z}};\alpha)\triangleq({\mathbf{A}}^{*}{\mathbf{A}}+\alpha)^{-1}({\mathbf{A}}^{*}{\mathbf{q}}+\alpha{\mathbf{z}}),$ (24)
$\text{prox}_{R}({\mathbf{z}};\eta)\triangleq\operatorname*{argmin}_{\mathbf{f}}\dfrac{1}{2\eta}\|{\mathbf{f}}-{\mathbf{z}}\|_{2}^{2}+R({\mathbf{f}}),$ (25)
@@
where ${\mathbf{A}}^{*}$ is the adjoint forward operator. The proximal operator in (25) can be interpreted as the denoiser of ${\mathbf{z}}$ with image prior $R(\cdot)$ and AWGN variance $\eta$. The key idea of the PnP is to employ a powerful image denoiser, like a pre-trained CNN, for (25). The convergence analysis of PnP is studied in the literature [107, 108, 109].
VII-F Radio interferometric imaging
@@
Radio interferometry is an imaging technique where we acquire samples of the Fourier transform of the image along curves [110, 111]. In radio interferometric imaging, aperture synthesis techniques are used to acquire specific Fourier measurements giving an incomplete coverage of the Fourier plane. The inverse problem consists of recovering the entire intensity image from Fourier measurements. In this case, the forward operator in a simplified form writes
${\mathbf{A}}=\mathbf{M}_{\text{RI}}\mathbf{F}$ (26)
@@
where $\mathbf{M}_{\text{RI}}$ is a Fourier mask indicating the Fourier coverage of the observations. We have omitted gridding and degridding operations in between other corrections from the forward operator for the sake of simplicity; a more detailed description of the operator can be found in  [110, §3.2].
@@
Figure 13: Training dataset used for image denoising experiment at resolution $512\times 512$.
VIII Futher analysis and additional experiments
@@
In this section we provide further analysis of LoFi architecture and additional experiments.
VIII-A High-dimensional image denoising
In this section, we train LoFi on 29900 training samples from the CelebA-HQ dataset at resolution $1024\times 1024$ for image denoising task ($\sigma=0.2$).
Figure 14 illustrates the reconstructed images for both in-distribution and OOD samples. This experiment shows that LoFi can achieve high-quality reconstructions for large images despite the simplicity of its architecture.
@@ VIII-A High-dimensional image denoising
Figure 14: LoFi reconstructions on in-distribution and OOD data for image denoising ($\sigma=0.2$) at resolution $1024\times 1024$.
VIII-B Equivariance to shifts and rotations
@@
In many imaging and pattern recognition tasks, it is desirable to work with functions that are invariant or equivariant to shifts and rotations. Building such invariances into CNNs, typically via group representation theory [112, 113] improves generalization both in and out of distribution [114, 115].
@@
It is easy to see that LoFi is shift equivariant by design. Indeed, the patch at position ${\mathbf{x}}+\Delta{\mathbf{x}}$ in an image shifted by $\Delta{\mathbf{x}}$ is exactly the same as the patch at position ${\mathbf{x}}$ in a non-shifted image; the $\text{NN}_{\theta}$ (MultiMLP) thus receives the same input. This stands in contrast to most mainstream multi-scale CNNs which are surprisingly sensitive even to 1-pixel shifts, although equivariance can be restored by a careful design of pooling and downsampling layers [116, 117].
@@
Rotation equivariance is precluded for square patches, except for angles that are multiples of $\frac{\pi}{2}$. With circular patches shown in Figure 3, when the image is rotated the $\text{NN}_{\theta}$ receives the same input only in a different order up to interpolation error. Rotation equivariance could be implemented by applying $\text{NN}_{\theta}$ to rotation invariant features [118, 119, 120, 121] extracted from the circular patch or by averaging the output over rotations. Even without this, however, LoFi is approximately rotation equivariant because natural images are approximately locally rotation invariant and the patch geometry is initialized as circular.
@@
Now, we assess LoFi performance on the image denoising task when the input image is shifted or rotated. As shown in Figure 15, we horizontally shifted the input noisy image for 15 pixels. As expected, LoFi’s reconstruction is shifted for 15 pixels and exhibits no degradation in the reconstruction quality. This experiment confirms that LoFi is a truly shift equivariant network.
@@
Figure 16 demonstrates the LoFi performance when the input noisy image undergoes a 90-degree rotation. As expected, we observe minimal degradation in reconstruction quality which shows LoFi’s strong robustness under rotation. Note that, we did not use any data augmentation to make the model robust under translation and rotation during training.
@@
Figure 15: Shift equivariance analysis for LoFi architecture; LoFi exhibits no degradation in reconstruction quality when the input noisy image is shifted for 15 pixels.
@@
Figure 16: Rotation equivariance analysis for LoFi architecture; LoFi exhibits minimal degradation in reconstruction quality when the input noisy image undergoes a 90-degree rotation.
Figure 17: The learned filter ${\mathbf{H}}$ in (15).